OpenStudy (anonymous):

Write an equation of a line in point-slope form that has a slope of -2 and passes through (5, -1). A. y + 1 = -2(x – 5) B. y – 1 = -2(x – 5) C. y – 5 = -2(x + 1) D. y -5 = -2(x – 1)

OpenStudy (anonymous):

Equation of line in slope -point form is given as \[y-y_1=m(x-x_1)\] Where slope m= -2 x1= 5 and y1= -1 Substituting these values in the above formula we have \[y-(-1)=-2(x-5) \rightarrow y+1=-2(x-5)\] Hence option A. y + 1 = -2(x – 5) is the correct answer. @mikebyy